WE HAVE 5ML OF 80% ACID WE NEED TO DILUTE IT WITH 2% ACID TO GET 10%. HOW MUCH 2%DO WE NEED?
MARILYN WANTS TO COMBINE A 30% ALCOHOL SOLUTION WITH 50% ALCOHOL SOLUTION TO GET 500MG OF AN 80% ALCOHOL SOLUTION PLEASES HELP HER ACCOMPLISH THIS TASK?

Found 2 solutions by stanbon, Paul:
Answer by stanbon(57984)

#1. Follow the acid.
Start with 80%(5)
Add 2%(x); x because the amount of fluid added is unknown
End up with: 10%(5+x)
EQUATION:
80%(5)+2%(x)=10%(5+x)
Multiply through by 100 to get:
80(5)+2x=10(5+x)
400+2x=50+10x
8x=350
x= 43.75ml (amount of 2% fluid you need to add)
Cheers,
Stan H.

Answer by Paul(988)

1.
80(5)+2x=10(x+5)
400+2x=10x+50
-8x=-350
x=43.75
Hence, about 43.75ml of sloution is needd.

2.80(500)-30x=50(500+x)
40000-30x=50x+25000
-80x=-15000
x=187.5
500-187.5 = 312.5

Hence, about 187.5mg of 30% is needed and about 312.5mg of 50% is needed.
Paul.